OpenStudy (anonymous):

Is 4/6 and 10/15 proportional?

OpenStudy (anonymous):

write each fraction in lowest terms. if they will be equal then proportional, if not - then not proportional.

OpenStudy (anonymous):

If the numbers are even?

OpenStudy (anonymous):

@ksanka

OpenStudy (anonymous):

no it is not necessary. the numbers 4 and 6 have the GCF and numbers 10 and 15 also have GCF. For 4 and 6 GCF is 2 and for 10 and 15 it is 5. Now divide 4 and 6 by 2 and 10 and 15 by 5.

OpenStudy (anonymous):

4/6=2/3 10/15=2/3 since in lowest terms the fractions are equal, they are proportional.

OpenStudy (uniqueartist1):

yes

OpenStudy (anonymous):

yes they are equal